Delta corporation has common stock of $50 million, preferred stock of $20 million, and total liabilities of $30 million. its cost of equity is 12 percent, the cost of preferred stock is 9percent, and the before-tax cost of debt is 16 percent. the tax rate is 40 percent. what is deltas WACC?
step by step solutions. no excel please
Answer #1

Weighted average cost of capital (WACC) of firm can be computed with following equation:

W ACC = we * ke + wy* kp + wd* kd

where,

w = weight

ke = cost of equity

kp = cost of preferred stock

kd = Post tax cost of debt.

Common stock = $50 million

Preferred stock = $20 million

Debt(liabilities) = $30 million

Thus,

Total Capital = 50+20+30 = $100 million

we = 50/100 = 0.5

wp = 20/100 = 0.2

wd = 30/100 = 0.3

Cost of equity (ke) = 12% = 0.12

Cost of preferred stock (kp) = 9% = 0.09

Cost of Debt = 16%

Tax rate = 40%

Post tax cost of debt (kd) = 16%(1-40%) = 9.6% = 0.096

Putting the values

W ACC = 0.5 * 0.12 +0.2 * 0.09 +0.3 * 0.096

WACC = 0.1068

WACC = 10.68%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.